9lb of bleached flour which costs $8/lb were combined with 3lb of unbleached flour which costs $4/lb. Find the cost per lb of the mixture
Answer by checkley77(12844)

9*8+3*4=12x
72+12=12x
84=12x
x=84/12
x=$7 per lb. is the price of the mix.
Proof:
9*8+3*4=12*7
72+12=84
84=84